The pollution levels for two firms, A and B, are Qa and Qb, respectively, measured in tons per year. Say profit for A as a function of pollution is 25Qa – Qa2 / 2 so that the marginal benefits of pollution function is MBA = 25-Qa. Similarly, profits for B = 40Qb – 2Qb2 so that MBB= 40-4Qb.  How much is each firm going to produce unregulated in the market?

Select one:

a. Qa = Qb = 10 tons

b. 35 tons total

c. Qa = 25 tons; Qb = 10 tons

d. Qa = 50 tons, Qb = 20 tons
Answer #1

Ans.- (C)

They will produce until marginal benefit for each firm is 0.

MBA = 25 -Qa

0 = 25- Qa

Qa = 25

MBB = 40-4Qb

0 = 40-4Qb

Qb= 10